OpenStudy (anonymous):

r^2+13r-30=0 what do you do after (r+15)(r-2)=0 ? thanks

OpenStudy (lalaly):

r=-15 , r=2

OpenStudy (anonymous):

r+15=0 or r-2=0 since multiplication

OpenStudy (anonymous):

sorry don't understand step by step with this equation please thanks

OpenStudy (anonymous):

r^2+13r-30=0,(r+15)(r-2)=0, r+15=0 or r-2=0, r=-15 or r=2